solve by the substitution method
4m+n=17
m-2n=20

Answer by rfer(16322)

m-2n=20
m=2n+20
-------------
4m+n=17
4(2n+20)+n=17
8n+80+n=17
9n=-63
n=-7
--------------
m-2(-7)=20
m+14=20
m=6
---------------
4(6)-7=17